OpenStudy (anonymous):

The cost in dollars of producing x units of a product is given by C(x) = \frac{7 x^2 - 27 x + 12}{3 \sqrt{x}} for x \ge 0. Determine the value of x when the marginal cost is 0

OpenStudy (anonymous):

Wow...that is a mess.

OpenStudy (anonymous):

\[\frac{7 x^2 - 27 x + 12}{3 \sqrt{x}} for x \ge 0.\]This is it fixed.

OpenStudy (anonymous):

I believe it's, anyways.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Set the derivative to zero and find the value of x at that point?

OpenStudy (anonymous):

Just wondering, that's undefined when x = 0, right? Did you mean x > 0?

OpenStudy (anonymous):

x is more than or = to zero

OpenStudy (anonymous):

@rukh - I can help you with the apartment problem too.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

That one, the occupancy equation is -8x + 960 (since at $960, you get zero occupancy). Rent revenue = -8x^2 + 960x. Differentiate and set to zero, you get x = 60 units. At x = 60 units, price = -480 + 960 = $480.

OpenStudy (anonymous):

Revenue is maximized at price $480, which results in occupancy of 60 units.

OpenStudy (anonymous):

thanks, that was correct...60 units

OpenStudy (anonymous):

now what abt this on with margianl cos?

OpenStudy (anonymous):

I am not too good with derivatives. Forgot long ago. But, to get marginal cost = 0, you are finding the tangent of the curve with slope = 0. That turns out to be the derivative of the cost function set to zero.

OpenStudy (anonymous):

but then i get 2 values for x

OpenStudy (anonymous):

and i dont think either one is the answer

OpenStudy (anonymous):

What are the values?

OpenStudy (anonymous):

GT, im curious how u know if derivative of c(x) gives u "marginal cost" ?

OpenStudy (anonymous):

the derivative of cost is the marginal cost

OpenStudy (anonymous):

Marginal cost = increase in cost to produce extra units of the product. That cost is zero where the "curve" flattens. That means, tangent to curve has slope zero.

OpenStudy (anonymous):

so how do i calculate?

OpenStudy (anonymous):

@saifoo.khan Can you help with derivative of this above function? Need values of x where that derivative is zero.

OpenStudy (anonymous):

i c, gota review my english market vocabs =(

OpenStudy (anonymous):

That's when:\[7x^{2} - 9x - 4 = 0\]as\[\frac{dC}{dx} = \frac{7x^{2} - 9x - 4}{2x^{3/2}}\]

OpenStudy (anonymous):

how is dC/dx that?

OpenStudy (anonymous):

ok so whats x?

OpenStudy (anonymous):

I think you need to treat it as f(x) * g(x) where 7x^2-9x-4 is f(x) and 1/3x^1/2 is g(x) and apply d/dx(f(x)*g(x)) = f(x)*dg(x)/dx + g(x)*df(x)/dx.

OpenStudy (anonymous):

I double checked my answer with the Wolf. Click show steps for derivation: http://www.wolframalpha.com/input/?i=d%28%287*x%5E2+-+27*x+%2B+12%29%2F%283*sqrt%28x%29%29%29%2Fdx

OpenStudy (saifoo.khan):

@GT , sorriieee. i just saw the notif. do you need help now as well?

OpenStudy (anonymous):

hey guys i have 4 mins to solve this problem before my hw expires :(

OpenStudy (anonymous):

Well, yeah.

OpenStudy (anonymous):

The derivative of that function.

OpenStudy (saifoo.khan):

wolf shows the steps as well, that will help.

OpenStudy (anonymous):

@bmp - if it is when 7x2−9x−4=0, then quadratic roots of that are: (9 +/- SQRT(81+112))/14 (9 +/- SQRT(193))/14 Since, x can't be negative, it has to be: (9 + SQRT(193))/14. Right.

OpenStudy (anonymous):

1.635??

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

@GT I got the same value for x, at least. :-)

OpenStudy (anonymous):

:)